OpenStudy (anonymous):

4x-2y=-2 3x+y=-4 substitution or elimination to solve

OpenStudy (anonymous):

You can solve for y in the second equation and then substitute into the first one to solve for x

OpenStudy (anonymous):

elimination of x or y

OpenStudy (anonymous):

\[y = -4-3x\] then substitute in the first equation for y

OpenStudy (anonymous):

\[4x-2(-4-3x)=-2\]

OpenStudy (anonymous):

.

OpenStudy (anonymous):

then distribute \[4x+8+6x=-2\] then combine like terms \[10x+8=-2\]

OpenStudy (anonymous):

solve for x \[10x=-10\] \[x=-1\]

OpenStudy (anonymous):

Plug that value into one of your equations and you get \[y = -1\]

OpenStudy (anonymous):

mul eq1 by 3 and mul eq2 by 4 then subtract (eq1 -eq2)

OpenStudy (anonymous):

That works also

OpenStudy (anonymous):

ok Thank you! I'm going to do one on my own to see if I completely understand..

OpenStudy (anonymous):

y=2x-7 x+4y=-1 my answer is (3,-1)

OpenStudy (anonymous):

Not sure those values are correct try plugging them into the second equation

OpenStudy (anonymous):

ok, yes it went in. Thank you

OpenStudy (anonymous):

oh you are right so sorry

OpenStudy (anonymous):

my mistake very good!

OpenStudy (anonymous):

I just hope I remeber when it comes to finals this week! UGH! night thank you